# Identifying gap-closings in open non-Hermitian systems by Biorthogonal Polarization

### v2

@@ I Introduction
1d
2d
Eigenvalues: $B_{2}=\mathbb{Z}$
Eigenvalues: $(a,b)\in(\mathbb{Z},\mathbb{Z})$, both even
@@ II Identifying phase transitions for open boundaries
in one and two spatial dimensions. This form of $H$ has a chiral symmetry such that the eigenvalues come in pairs of ±$\pm E$. This means an EP can appear only at $E=0$. We consider an open system with OBC imposed along one of the directions. Defining $H({\mathbf{k}_{\perp}},\beta)=\mathbf{d}({\mathbf{k}_{\perp}},\beta)\cdot\boldsymbol{\sigma}$, where $\beta$-substitution corresponds to the momentum-component along which OBC is imposed, and ${\mathbf{k}_{\perp}}$ denotes the momentum directions along which periodicity is retained. According to the prescription explained above, the characteristic polynomial for $\beta$ takes the form:
@@
This section is devoted to the study of non-Hermitian generalizations of the 1d SSH model [34, 35]. A tight-binding model for the non-Hermitian SSH chain is obtained by setting [18, 36, 37]
@@
in Eq. (3). The corresponding lattice Hamiltonian is given by
@@
Figure 2: Absolute values of the eigenvalue spectrum for the Hamiltonian in Eq. (III), with the parameter values (a) $t_{2}=1$, $t_{3}=0$, $\gamma=3$, $\Delta=0$; (b) $t_{2}=1$, $t_{3}=0$, $\gamma=3$, $\Delta=1$; (c) $t_{2}=1$, $t_{3}=1/5$, $\gamma=4/3$, $\Delta=0$; (d) $t_{2}=1$, $t_{3}=1/5$, $\gamma=4/3$, $\Delta=1$. The spectra for the bulk states, corresponding to PBC and OBC cases, have been shown in gray and blue, respectively. The edge mode energies have been indicated with green. The biorthogonal polarization $P$ (depicted via the dashed red lines) jumps precisely by unity when the edge enters or leaves the bulk spectrum (blue), and these jumps have no correspondence with the location of the EPs of the periodic spectrum. We have also shown the total non-Bloch band theory winding $\nu_{tot}$ (demarcated by cyan curves) for the two bands. In (a) and (b), $\nu_{tot}$ coincides with $P$. But in (c) and (d), $\nu_{tot}$ deviates from a quantized value. Nevertheless, it still shows jumps at the phase transitions points captured by $P$. Denoting the number of unit cells taken for plotting the OBC spectra and computing $P$ by $N_{\textrm{OBC}}$ and $N_{P}$, respectively, we have used $\{N_{\textrm{OBC}},N_{P}\}$ equal to (a) $\{100,3500\}$; (b) $\{100,3000\}$; (c) $\{90,90\}$; and (d) $\{90,90\}$.
@@
The nonzero $t_{3}$ case corresponds to longer-ranged hoppings (in the sense of the $A$-$B$ sublattice-site-arrangements of Fig. 1 of Ref. [18] and Fig. 2 of Ref. [37]) and, hence, the open chain does not admit analytical expressions for the energy eigenvalues, eigenvectors, or BP. In fact, the polynomials in Eqs. (6) and (7) are fourth order in $\beta$, and the admissible solutions are the ones when the roots satisfy $|\beta_{2}|=|\beta_{3}|$ (after ordering the $\beta_{j}$’s according to their magnitude). This gives us the GBZ, and we find that it forms a closed loop in the complex $\beta$-plane where the magnitude of $\beta$ also changes. In other words, unlike the $t_{3}=0$ case, the GBZ is no longer a circle. The Hamiltonian $H(\beta)$, with the admissble values of $\beta$ fed in, give us the energy and eigenvctors of the open system. Alternatively, we can of course determine all these features by considering the Hamiltonian of the open chain in the real space, with $N$ unit cells. In Fig. 2(c) and (d), we have computed the spectrum and $P$ by considering the real-space Hamiltonian of length $N=45$, i.e., with a broken unit cell at $N=46$. On the other hand, the non-Bloch or generalized winding vector has been calculated by computing the GBZ.
@@ IV.1.1 Model for $t_{3}=0$ admitting closed-form analytical solutions
Figure 3: Absolute values of the energy spectrum for the Hamiltonian ${\mathbf{d}}({\mathbf{k}})\cdot\boldsymbol{\sigma}$, with the components of ${\mathbf{d}}$ given by Eq. (IV), after imposing OBC along the $x$-axis. The chosen parameter values are (a) $t_{1}=\delta=\Delta=1$, $t_{3}=0$, $\gamma=3$ (cf. the model of Sec. IV.1.1 admitting closed-form analytical solutions) and (b) $t_{1}=\delta=2$, $t_{3}=1/2$, $\gamma=4/5$, $\Delta=1/4$. The spectra for the bulk states, corresponding to the PBC and OBC cases, have been shown in gray and blue, respectively. The edge mode energies have been indicated with green. The biorthogonal polarization $P$ (depicted via the dashed red lines) jumps precisely by unity when the edge enters or leaves the bulk spectrum (blue), and these jumps have no correspondence with the location of the EPs of the periodic spectrum. We have also shown the total non-Bloch band theory winding $\nu_{tot}$ (demarcated by cyan curves) for the two bands. In (a), $\nu_{tot}$ coincides with $P$. But in (b), $\nu_{tot}$ deviates from a quantized value, as for non-chiral systems, such a topological number does not always take quantized values. Nevertheless, it still shows jumps at the phase transitions points captured by $P$. Denoting the number of unit cells taken for plotting the OBC spectra and computing $P$ by $N_{\textrm{OBC}}$ and $N_{P}$, respectively, we have used $\{N_{\textrm{OBC}},N_{P}\}$ equal to (a) $\{80,3000\}$ and (b) $\{100,1000\}$.
@@ IV.2 OBC along the $y$-axis
Figure 5: Absolute values of the energy spectrum for the Hamiltonian ${\mathbf{d}}({\mathbf{k}})\cdot\boldsymbol{\sigma}$, with the components of ${\mathbf{d}}$ given by (a) Eq. (39) and (b) Eq. (42), after imposing OBC along the $y$-axis. For both the subfigures, the parameter values have been set to $t_{1}=1$, $\delta=1.75$, $t_{3}=0$, $\gamma=2/5$, and $\Delta=0.1$. The spectra for the bulk states, corresponding to the PBC and OBC cases, have been shown in gray and blue, respectively. The edge mode energies have been indicated with green. The biorthogonal polarization $P$ (depicted via the dashed red lines) jumps precisely by unity when the edge enters or leaves the bulk spectrum (blue), and these jumps have no correspondence with the location of the EPs of the periodic spectrum. We have also shown the total non-Bloch band theory winding $\nu_{tot}$ (demarcated by cyan curves) for the two bands. In both cases, $\nu_{tot}$ does not show any jump at the phase transitions points captured by $P$. Denoting the number of unit cells taken for plotting the OBC spectra and computing $P$ by $N_{\textrm{OBC}}$ and $N_{P}$, respectively, we have used $\{N_{\textrm{OBC}},N_{P}\}=\{100,1200\}$ for both the subfigures.

### v3

@@ I Introduction
1d 2d
Eigenvalues: $B_{2}=\mathbb{Z}$ Eigenvalues: $(a,b)\in(\mathbb{Z},\mathbb{Z})$, both even
@@ II Identifying phase transitions for open boundaries
in one and two spatial dimensions. A Hamiltonian of the form ${\mathbf{d}}({\mathbf{k}})\cdot\boldsymbol{\sigma}$ always has the eigenvalues in pairs of ±$\pm E$. This means an EP can appear only at $E=0$. We consider an open system with OBC imposed along one of the directions. Defining $H({\mathbf{k}_{\perp}},\beta)=\mathbf{d}({\mathbf{k}_{\perp}},\beta)\cdot\boldsymbol{\sigma}$, where $\beta$-substitution corresponds to the momentum-component along which OBC is imposed, and ${\mathbf{k}_{\perp}}$ denotes the momentum directions along which periodicity is retained. According to the prescription explained above, the characteristic polynomial for $\beta$ takes the form:
@@
This section is devoted to the study of non-Hermitian generalizations of the 1d SSH model [34, 35] and Rice-Mele chains [38, 44]. A tight-binding model for such 2d chains is obtained by setting [18, 36, 37]
@@
in Eq. (3). The SSH version corresponds to $\Delta=0$. The corresponding lattice Hamiltonian is given by
@@
Figure 2: Absolute values of the eigenvalue spectrum for the Hamiltonian in Eq. (III), with the parameter values (a) $t_{2}=1$, $t_{3}=0$, $\gamma=3$, $\Delta=0$; (b) $t_{2}=1$, $t_{3}=0$, $\gamma=3$, $\Delta=1$; (c) $t_{2}=1$, $t_{3}=1/5$, $\gamma=4/3$, $\Delta=0$; (d) $t_{2}=1$, $t_{3}=1/5$, $\gamma=4/3$, $\Delta=1$. The spectra for the bulk states, corresponding to PBC and OBC cases, have been shown in gray and blue, respectively. The edge mode energies have been indicated with green. The biorthogonal polarization $P$ (depicted via the dashed red lines) jumps precisely by unity when the edge enters or leaves the bulk spectrum (blue), and these jumps have no correspondence with the location of the EPs of the periodic spectrum. We have also shown the total non-Bloch band theory winding $\nu_{tot}$ (demarcated by cyan curves) for the two bands. In (a) and (b), $\nu_{tot}$ coincides with $P$. But in (c) and (d), $\nu_{tot}$ deviates from a quantized value. Nevertheless, it still shows jumps at the phase transitions points captured by $P$. Denoting the number of unit cells taken for plotting the OBC spectra and computing $P$ by $N_{\textrm{OBC}}$ and $N_{P}$, respectively, we have used $\{N_{\textrm{OBC}},N_{P}\}$ equal to (a) $\{100,3500\}$; (b) $\{100,3000\}$; (c) $\{45,45\}$; and (d) $\{45,45\}$.
@@
The nonzero $t_{3}$ case corresponds to longer-ranged hoppings (in the sense of the $A$-$B$ sublattice-site-arrangements of Fig. 1 of Ref. [18] and Fig. 2 of Ref. [37]) and, hence, the open chain does not admit analytical expressions for the energy eigenvalues, eigenvectors, or BP. In fact, the polynomials in Eqs. (6) and (7) are fourth order in $\beta$, and the admissible solutions are the ones when the roots satisfy $|\beta_{2}|=|\beta_{3}|$ (after ordering the $\beta_{j}$’s according to their magnitude). This gives us the GBZ, and we find that it forms a closed loop in the complex $\beta$-plane where the magnitude of $\beta$ also changes. In other words, unlike the $t_{3}=0$ case, the GBZ is no longer a circle. The Hamiltonian $H(\beta)$, with the admissble values of $\beta$ fed in, give us the energy and eigenvctors of the open system. Alternatively, we can of course determine all these features by considering the Hamiltonian of the open chain in the real space, with $N$ unit cells. In Fig. 2(c) and (d), we have computed the spectrum and $P$ by considering the real-space Hamiltonian of length $N=45$. On the other hand, the non-Bloch or generalized winding vector has been calculated by computing the GBZ.
@@ IV.1.1 Model for $t_{3}=0$ admitting closed-form analytical solutions
Figure 3: Absolute values of the energy spectrum for the Hamiltonian ${\mathbf{d}}({\mathbf{k}})\cdot\boldsymbol{\sigma}$, with the components of ${\mathbf{d}}$ given by Eq. (IV), after imposing OBC along the $x$-axis. The chosen parameter values are (a) $t_{1}=\delta=\Delta=1$, $t_{3}=0$, $\gamma=3$ (cf. the model of Sec. IV.1.1 admitting closed-form analytical solutions) and (b) $t_{1}=\delta=2$, $t_{3}=1/2$, $\gamma=4/5$, $\Delta=1/4$. The spectra for the bulk states, corresponding to the PBC and OBC cases, have been shown in gray and blue, respectively. The edge mode energies have been indicated with green. The biorthogonal polarization $P$ (depicted via the dashed red lines) jumps precisely by unity when the edge enters or leaves the bulk spectrum (blue), and these jumps have no correspondence with the location of the EPs of the periodic spectrum. We have also shown the total non-Bloch band theory winding $\nu_{tot}$ (demarcated by cyan curves) for the two bands. In (a), $\nu_{tot}$ coincides with $P$. But in (b), $\nu_{tot}$ deviates from a quantized value, as for non-chiral systems, such a topological number does not always take quantized values. Nevertheless, it still shows jumps at the phase transitions points captured by $P$. Denoting the number of unit cells taken for plotting the OBC spectra and computing $P$ by $N_{\textrm{OBC}}$ and $N_{P}$, respectively, we have used $\{N_{\textrm{OBC}},N_{P}\}$ equal to (a) $\{80,3000\}$ and (b) $\{50,500\}$.
@@ IV.2 OBC along the $y$-axis
Figure 5: Absolute values of the energy spectrum for the Hamiltonian ${\mathbf{d}}({\mathbf{k}})\cdot\boldsymbol{\sigma}$, with the components of ${\mathbf{d}}$ given by (a) Eq. (28) and (b) Eq. (31), after imposing OBC along the $y$-axis. For both the subfigures, the parameter values have been set to $t_{1}=1$, $\delta=1.75$, $t_{3}=0$, $\gamma=2/5$, and $\Delta=0.1$. The spectra for the bulk states, corresponding to the PBC and OBC cases, have been shown in gray and blue, respectively. The edge mode energies have been indicated with green. The biorthogonal polarization $P$ (depicted via the dashed red lines) jumps precisely by unity when the edge enters or leaves the bulk spectrum (blue), and these jumps have no correspondence with the location of the EPs of the periodic spectrum. We have also shown the total non-Bloch band theory winding $\nu_{tot}$ (demarcated by cyan curves) for the two bands. In both cases, $\nu_{tot}$ does not show any jump at the phase transitions points captured by $P$. Denoting the number of unit cells taken for plotting the OBC spectra and computing $P$ by $N_{\textrm{OBC}}$ and $N_{P}$, respectively, we have used $\{N_{\textrm{OBC}},N_{P}\}=\{50,600\}$ for both the subfigures.